# Validity of relativistic hydrodynamics beyond local equilibrium

### v1

@@ B.3 Hydrodynamics
$\partial_{\tau}\pi_{1,1}+\frac{\pi_{1,1}}{\tau_{R}}=(c_{s}^{2}-3)\frac{\pi_{1,1}}{\tau}+c_{s}^{2}\frac{(\epsilon+P)}{\tau}+\frac{\kappa}{\tau}+\frac{\rho^{eq}_{1,2}(\tau)}{\tau}-3\frac{P}{\tau}+\frac{\pi_{1,2}}{\tau}$ (155)
Evolution of $\pi_{-1,0}$
$\mathbf{\hat{D}}\pi_{-1,0}=-\frac{\pi_{-1,0}}{\tau_{R}}-\mathbf{\hat{D}}\rho^{eq}_{-1,0}(\tau)$ (156)
We note that $m^{2}\rho^{eq}_{-1,0}(\tau)=\langle m^{2}\rangle^{eq}=\epsilon-3P$ and we get,
$\mathbf{\hat{D}}\rho^{eq}_{-1,0}(\tau)$ $=\frac{1}{m^{2}}\mathbf{\hat{D}}(\epsilon-3P)$
$=\frac{1}{m^{2}}\mathbf{\hat{D}}\epsilon-3\mathbf{\hat{D}}P$
$=-3\frac{1}{m^{2}}\mathbf{\hat{D}}P\,$, (157)
where we have used the nergy conservation relation, $\mathbf{\hat{D}}\epsilon=0$. then the evolution equation for $\pi_{-1,0}$ reads,
$\mathbf{\hat{D}}\pi_{-1,0}=-\frac{\pi_{-1,0}}{\tau_{R}}+\frac{3}{m^{2}}\mathbf{\hat{D}}P$ (158)
Now,
$\mathbf{\hat{D}}\pi_{-1,0}$ $=\partial_{\tau}\pi_{1,1}+(2\cdot 0+1)\frac{\pi_{-1,0}}{\tau}-(2\cdot 0+1)\frac{\pi_{-1,1}}{\tau}$
$=\partial_{\tau}\pi_{-1,0}+\frac{\pi_{-1,0}}{\tau}-\frac{\pi_{-1,1}}{\tau}\,.$ (159)
Combining we have
$\partial_{\tau}\pi_{-1,0}+\frac{\pi_{-1,0}}{\tau_{R}}=-\frac{\pi_{-1,0}}{\tau}+\frac{3}{m^{2}}\left(-c_{s}^{2}\frac{(\epsilon+P+\pi_{1,1})}{\tau}-\frac{\kappa}{\tau}+\frac{3P}{\tau}-\frac{\rho^{eq}_{1,2}(\tau)}{\tau}\right)+\frac{\pi_{-1,1}}{\tau}$ (160)
Evolution of $\Pi$
Multiplying both sides of (158) by $\frac{m^{2}}{3}$, we get the evolution equation for $\Pi$
$\mathbf{\hat{D}}\Pi=-\frac{\Pi}{\tau_{R}}-\mathbf{\hat{D}}P$ (161)
From (160),
$\partial_{\tau}\Pi+\frac{\Pi}{\tau_{R}}$ $=-\frac{\Pi}{\tau}-\left(-c_{s}^{2}\frac{(\epsilon+P+\pi_{1,1})}{\tau}-\frac{\kappa}{\tau}+\frac{3P}{\tau}-\frac{\rho^{eq}_{1,2}(\tau)}{\tau}\right)-\frac{m^{2}}{3}\frac{\pi_{-1,1}}{\tau}$ (162)
$=-\frac{\Pi}{\tau}+c_{s}^{2}\frac{(\epsilon+P+\Pi-\pi)}{\tau}+\frac{\kappa}{\tau}-\frac{3P}{\tau}+\frac{\rho^{eq}_{1,2}(\tau)}{\tau}-\frac{m^{2}}{3}\frac{\pi_{-1,1}}{\tau}$ (163)
$=(1-c_{s}^{2})\frac{\Pi}{\tau}-\frac{\pi}{\tau}+c_{s}^{2}\frac{(\epsilon+P)}{\tau}+\frac{\kappa}{\tau}-\frac{3P}{\tau}+\frac{\rho^{eq}_{1,2}(\tau)}{\tau}-\frac{m^{2}}{3}\frac{\pi_{-1,1}}{\tau}$ (164)
Evolution of $\pi$
$\mathbf{\hat{D}}\pi_{1,1}=-\frac{\pi_{1,1}}{\tau_{R}}-\mathbf{\hat{D}}P$ (165)
Remembering $\pi_{1,1}=\Pi-\pi$,
$\mathbf{\hat{D}}(\Pi-\pi)=-\frac{\Pi-\pi}{\tau_{R}}-\mathbf{\hat{D}}P$ (166)
and
$\mathbf{\hat{D}}\Pi=-\frac{\Pi}{\tau_{R}}-\mathbf{\hat{D}}P$ (167)
Subtracting, we get
$\mathbf{\hat{D}}\pi=-\frac{\pi}{\tau_{R}}$ (168)
$\partial_{\tau}\pi_{1,1}+\frac{\pi_{1,1}}{\tau_{R}}=(c_{s}^{2}-3)\frac{\pi_{1,1}}{\tau}+c_{s}^{2}\frac{(\epsilon+P)}{\tau}+\frac{\kappa}{\tau}+\frac{\rho^{eq}_{1,2}(\tau)}{\tau}-3\frac{P}{\tau}+\frac{\pi_{1,2}}{\tau}$ (169)
$\partial_{\tau}(\pi-\Pi)+\frac{(\pi-\Pi)}{\tau_{R}}=(c_{s}^{2}-3)\frac{(\pi-\Pi)}{\tau}-c_{s}^{2}\frac{(\epsilon+P)}{\tau}-\frac{\kappa}{\tau}-\frac{\rho^{eq}_{1,2}(\tau)}{\tau}+3\frac{P}{\tau}-\frac{\pi_{1,2}}{\tau}$ (170)
rewriting,
$\partial_{\tau}\pi+\frac{\pi}{\tau_{R}}$ $=\partial_{\tau}\Pi+\frac{\Pi}{\tau_{R}}+(c_{s}^{2}-3)\frac{(\pi-\Pi)}{\tau}-c_{s}^{2}\frac{(\epsilon+P)}{\tau}-\frac{\kappa}{\tau}-\frac{\rho^{eq}_{1,2}(\tau)}{\tau}+3\frac{P}{\tau}-\frac{\pi_{1,2}}{\tau}$ (171)
$=(1-c_{s}^{2})\frac{\Pi}{\tau}-\frac{\pi}{\tau}+c_{s}^{2}\frac{(\epsilon+P)}{\tau}+\frac{\kappa}{\tau}-\frac{3P}{\tau}+\frac{\rho^{eq}_{1,2}(\tau)}{\tau}-\frac{m^{2}}{3}\frac{\pi_{-1,1}}{\tau}$ (172)
$+(c_{s}^{2}-3)\frac{(\pi-\Pi)}{\tau}-c_{s}^{2}\frac{(\epsilon+P)}{\tau}-\frac{\kappa}{\tau}-\frac{\rho^{eq}_{1,2}(\tau)}{\tau}+3\frac{P}{\tau}-\frac{\pi_{1,2}}{\tau}$ (173)
$=(1-c_{s}^{2})\frac{\Pi}{\tau}-\frac{\pi}{\tau}+(c_{s}^{2}-3)\frac{(\pi-\Pi)}{\tau}-\frac{m^{2}}{3}\frac{\pi_{-1,1}}{\tau}-\frac{\pi_{1,2}}{\tau}$ (174)
$=-(4-c_{s}^{2})\frac{\pi}{\tau}+2(2-c_{s}^{2})\frac{\Pi}{\tau}-\frac{m^{2}}{3}\frac{\pi_{-1,1}}{\tau}-\frac{\pi_{1,2}}{\tau}$ (175)
We have the finial expression
$\partial_{\tau}\pi+\frac{\pi}{\tau_{R}}=-(4-c_{s}^{2})\frac{\pi}{\tau}+2(2-c_{s}^{2})\frac{\Pi}{\tau}-\frac{m^{2}}{3}\frac{\pi_{-1,1}}{\tau}-\frac{\pi_{1,2}}{\tau}$ (176)
Typical structure of second order hydrodynamic equations
$\dot{\Pi}+\frac{\Pi}{\tau_{\Pi}}$ $=-\beta_{\Pi}\theta-\delta_{\Pi\Pi}\theta~\Pi+\lambda_{\Pi\pi}\sigma_{\mu\nu}\pi^{\mu\nu}\,.$ (177)
$\dot{\pi}^{\mu\nu}+\frac{\pi^{\mu\nu}}{\tau_{\pi}}$ $=2\beta_{\pi}\sigma^{\mu\nu}-\delta_{\pi\pi}\theta~\pi^{\mu\nu}+\lambda_{\pi\Pi}\Pi\sigma^{\mu\nu}$ (178)
$+2\pi_{\gamma}^{\langle\mu}\omega^{\nu\rangle\gamma}-\tau_{\pi\pi}+2\pi_{\gamma}^{\langle\mu}\sigma^{\nu\rangle\gamma}\,.$ (179)
Appendix C 3+13+1D Calculations
@@
$\left[\partial_{\tau}+\hat{\textbf{L}}+\hat{\boldsymbol{\nu}}_{R}\right]\vec{\rho}=\hat{\boldsymbol{\nu}}_{R}\vec{\rho}^{~\text{eq}}\,.$ (180)
@@ C.1 Collision picture
$\left[\partial_{\tau}+\hat{\mathbf{L}}\right]=0$ (181)
@@
$\partial_{\tau}\hat{\mathcal{F}}_{\tau,\tau_{0}}=-\hat{\mathbf{L}}~\hat{\mathcal{F}}_{\tau,\tau_{0}}$ (182)
@@
$\int_{\tau_{0}}^{\tau}d\tau^{\prime}\mathcal{W}_{\tau,\tau^{\prime}}~\hat{\boldsymbol{\nu}}_{c}(\tau^{\prime})\vec{\rho}_{c}^{~\text{eq}}(\tau^{\prime})$ $=\int_{\tau_{0}}^{\tau}d\tau^{\prime}\partial_{\tau^{\prime}}\mathcal{W}_{\tau,\tau^{\prime}}\rho_{c}^{~\text{eq}}(\tau^{\prime})$ (188)
$=\mathcal{W}_{\tau,\tau^{\prime}}\rho_{c}^{~\text{eq}}(\tau^{\prime})\Bigg{|}_{\tau_{0}}^{\tau}-\int_{\tau_{0}}^{\tau}d\tau^{\prime}\mathcal{W}_{\tau,\tau^{\prime}}\partial_{\tau^{\prime}}\rho_{c}^{~\text{eq}}(\tau^{\prime})$ (189)
$=\mathcal{W}_{\tau,\tau^{\prime}}\rho_{c}^{~\text{eq}}(\tau^{\prime})\Bigg{|}_{\tau_{0}}^{\tau}-\int_{\tau_{0}}^{\tau}d\tau^{\prime}\mathcal{W}_{\tau,\tau^{\prime}}\hat{\boldsymbol{\nu}}_{c}(\tau^{\prime})\hat{\boldsymbol{\nu}}_{c}^{-1}(\tau^{\prime})\partial_{\tau^{\prime}}\rho_{c}^{~\text{eq}}(\tau^{\prime})$ (190)
$=\sum_{n=0}^{\infty}\mathcal{W}_{\tau,\tau^{\prime}}~\left[\hat{\boldsymbol{\nu}}_{c}^{-1}(\tau^{\prime})\partial_{\tau^{\prime}}\right]^{k}~\rho_{c}^{~\text{eq}}(\tau^{\prime})\Bigg{|}_{\tau_{0}}^{\tau}$ (191)
$=\hat{\mathcal{F}}^{-1}_{\tau,\tau_{0}}\sum_{n=0}^{\infty}\left[\hat{\boldsymbol{\nu}}^{-1}(\tau)\mathbf{\hat{D}}\right]^{k}~\rho^{~\text{eq}}(\tau)-\mathcal{W}_{\tau,\tau_{0}}~\sum_{n=0}^{\infty}\left[\hat{\boldsymbol{\nu}}^{-1}(\tau_{0})\mathbf{\hat{D}}\right]^{k}~\rho^{~\text{eq}}(\tau_{0})$ (192)
@@
$\vec{\rho}(\tau)=\hat{\mathcal{F}}_{\tau,\tau_{0}}\mathcal{W}_{\tau,\tau_{0}}\rho(\tau_{0})+\sum_{n=0}^{\infty}\left[\hat{\boldsymbol{\nu}}^{-1}(\tau)\mathbf{\hat{D}}\right]^{k}~\rho^{~\text{eq}}(\tau)-\hat{\mathcal{F}}_{\tau,\tau_{0}}\mathcal{W}_{\tau,\tau_{0}}~\sum_{n=0}^{\infty}\left[\hat{\boldsymbol{\nu}}^{-1}(\tau_{0})\mathbf{\hat{D}}\right]^{k}~\rho^{~\text{eq}}(\tau_{0})$ (193)
C.2 Free Streaming picture
@@
$\left[\partial_{\tau}+\hat{\mathbf{L}}_{S}\right]\vec{\rho}_{S}=\partial_{\tau}\vec{\rho}_{S}$ (199)
@@
$\hat{\mathcal{F}}_{\tau,\tau_{0}}=\mathcal{T}e^{-\int_{\tau_{0}}^{\tau}d\tau\hat{\textbf{L}}_{S}}$ (200)
@@
$\partial_{\tau}\hat{\mathcal{F}}_{\tau,\tau_{0}}=-\hat{\textbf{L}}_{S}~\hat{\mathcal{F}}_{\tau,\tau_{0}}$ (201)
@@
$\vec{\rho}^{G}(\tau)$ $=\int_{\tau_{0}}^{\tau}d\tau^{\prime}e^{-\xi^{\prime}}\hat{\mathcal{F}}_{\tau,\tau^{\prime}}\frac{\rho^{~\text{eq}}(\tau^{\prime})}{\tau_{R}}$ (205)
@@
$\vec{\rho}^{G}(\tau)$ $=\sum_{n=0}^{\infty}\frac{g_{n}(\tau,\tau_{0})}{n!}\sum_{k=0}^{n}h_{n,k}(\tau_{R},\mathbf{\hat{D}}\tau_{R})\left[-\tau_{R}\mathbf{\hat{D}}\right]^{k}\vec{\rho}^{~\text{eq}}$ (209)
$=\sum_{n=0}^{\infty}~\left(\sum_{nk=n}^{\infty}\frac{g_{k}(\tau,\tau_{0})}{n!}h_{k,n}(\tau_{R},\mathbf{\hat{D}}\tau_{R})\right)~\left[-\tau_{R}\mathbf{\hat{D}}\right]^{n}\vec{\rho}^{~\text{eq}}$ (210)

### v2

@@ C.2 Evolution equations
$\partial_{\tau}\pi_{1,1}+\frac{\pi_{1,1}}{\tau_{R}}=-(3-c_{s}^{2})\frac{\pi_{1,1}}{\tau}+\left[c_{s}^{2}(\varepsilon+P)+\kappa-3\left(1-\frac{1}{5}\frac{G_{4,2}(z)}{G_{4,1}(z)}\right)P\right]\frac{1}{\tau}+\frac{\pi_{1,2}}{\tau}$ (214)
Now from the equation (214),
$\partial_{\tau}(\pi-\Pi)+\frac{(\pi-\Pi)}{\tau_{R}}=-(3-c_{s}^{2})\frac{(\pi-\Pi)}{\tau}-\left[c_{s}^{2}(\varepsilon+P)+\kappa-3\left(1-\frac{1}{5}\frac{G_{4,2}(z)}{G_{4,1}(z)}\right)\right]\frac{1}{\tau}-\frac{\pi_{1,2}}{\tau}$ (215)
$\partial_{\tau}\pi+\frac{\pi}{\tau_{R}}$ $=\left(\partial_{\tau}\Pi+\frac{\Pi}{\tau_{R}}\right)$ (216)
$~~~~~~~~~~~~-(3-c_{s}^{2})\frac{(\pi-\Pi)}{\tau}-\left[c_{s}^{2}(\varepsilon+P)+\frac{\kappa}{\tau}-3\left(1-\frac{1}{5}\frac{G_{4,2}(z)}{G_{4,1}(z)}\right)P\right]\frac{1}{\tau}-\frac{\pi_{1,2}}{\tau}$
$=-\left[\frac{4}{3}-c_{s}^{2}\right]\frac{\Pi}{\tau}+\left[\frac{1}{3}-c_{s}^{2}\right]\frac{\pi}{\tau}-\left[\left(\frac{1}{3}-c_{s}^{2}\right)(\varepsilon+P)-\kappa-\frac{1}{3}\left(1-\frac{1}{3}\frac{G_{2,1}(z)}{G_{2,0}(z)}\right)(\varepsilon-3P)\right]\frac{1}{\tau}-\frac{m^{2}}{3}\frac{1}{\tau}\pi_{-1,1}$
$~~~~~~~~~~~~~~~-(3-c_{s}^{2})\frac{(\pi-\Pi)}{\tau}-\left[c_{s}^{2}(\varepsilon+P)+\kappa-3\left(1-\frac{1}{5}\frac{G_{4,2}(z)}{G_{4,1}(z)}\right)P\right]\frac{1}{\tau}-\frac{\pi_{1,2}}{\tau}$
$=-\frac{8}{3\tau}\pi+\frac{5}{3\tau}\Pi-\left[\frac{1}{3}(\varepsilon+P)-\frac{1}{3}\left(1-\frac{1}{3}\frac{G_{2,1}(z)}{G_{2,0}(z)}\right)(\varepsilon-3P)-3\left(1-\frac{1}{5}\frac{G_{4,2}(z)}{G_{4,1}(z)}\right)P\right]\frac{1}{\tau}-\frac{m^{2}}{3}\frac{\pi_{-1,1}}{\tau}-\frac{\pi_{1,2}}{\tau}$ (217)
Interestingly for the Conformal case, we have the relation,
$\partial_{\tau}\pi+\frac{\pi}{\tau_{R}}$ $=-(3-c_{s}^{2})\frac{\pi}{\tau}-\left[c_{s}^{2}(\varepsilon+P)+\frac{\kappa}{\tau}-3\left(1-\frac{1}{5}\frac{G_{4,2}(0)}{G_{4,1}(0)}\right)P\right]\frac{1}{\tau}-\frac{\pi_{1,2}}{\tau}$ (218)
$=-\frac{8}{3}\frac{\pi}{\tau}-\left[\frac{4}{3}P+0-\frac{3\cdot 4}{5}P\right]\frac{1}{\tau}-\frac{\pi_{1,2}}{\tau}$ (219)
$=-\frac{8}{3}\frac{\pi}{\tau}+\frac{4}{3}\left(\frac{4}{5}P\right)\frac{1}{\tau}-\frac{\pi_{1,2}}{\tau}$ (220)
C.3 Hydrodynamic equations
Collecting all the results we have for a massive system, the equations,
$\partial_{\tau}\pi+\frac{\pi}{\tau_{R}}$ $=-\frac{8}{3\tau}\pi+\frac{5}{3\tau}\Pi-\left[\frac{1}{3}(\varepsilon+P)-3\left(1-\frac{1}{5}\frac{G_{4,2}(z)}{G_{4,1}(z)}\right)P-\frac{1}{3}\left(1-\frac{1}{3}\frac{G_{2,1}(z)}{G_{2,0}(z)}\right)(\varepsilon-3P)\right]\frac{1}{\tau}-\frac{m^{2}}{3}\frac{\pi_{-1,1}}{\tau}-\frac{\pi_{1,2}}{\tau}$ (221)
$\partial_{\tau}\Pi+\frac{\Pi}{\tau_{R}}$ $=-\left[\frac{4}{3}-c_{s}^{2}\right]\frac{\Pi}{\tau}+\left[\frac{1}{3}-c_{s}^{2}\right]\pi-\left[\left(\frac{1}{3}-c_{s}^{2}\right)(\varepsilon+P)-\kappa-\frac{1}{3}\left(1-\frac{1}{3}\frac{G_{2,1}(z)}{G_{2,0}(z)}\right)(\varepsilon-3P)\right]\frac{1}{\tau}-\frac{m^{2}}{3}\frac{1}{\tau}\pi_{-1,1}\,.$ (222)
Comparison with DNMR
$\beta_{\Pi}=\left[\left(\frac{1}{3}-c_{s}^{2}\right)(\varepsilon+P)-\kappa-\frac{1}{3}\left(1-\frac{1}{3}\frac{G_{2,1}(z)}{G_{2,0}(z)}\right)(\varepsilon-3P)\right]$ (223)
For a conformal system we get the equation for sheer evolution,
$\partial_{\tau}\pi+\frac{\pi}{\tau_{R}}=-\frac{8}{3}\frac{\pi}{\tau}-\frac{4}{3}\left(\frac{4}{5}P\right)\frac{1}{\tau}-\frac{\pi_{1,2}}{\tau}$ (224)
Appendix D 3+13+1D Calculations
@@ C.3 Hydrodynamic equations
$\left[\partial_{\tau}+\hat{\textbf{F}}+\hat{\boldsymbol{\nu}}_{R}\right]\vec{\rho}=\hat{\boldsymbol{\nu}}_{R}\vec{\rho}^{~\text{eq}}\,.$ (225)
@@ D.1 Collision picture
$\left[\partial_{\tau}+\hat{\mathbf{F}}\right]=0$ (226)
@@
$\partial_{\tau}\hat{\mathcal{F}}_{\tau,\tau_{0}}=-\hat{\mathbf{F}}~\hat{\mathcal{F}}_{\tau,\tau_{0}}$ (227)
@@
$\int_{\tau_{0}}^{\tau}d\tau^{\prime}\mathcal{W}_{\tau,\tau^{\prime}}~\hat{\boldsymbol{\nu}}_{c}(\tau^{\prime})\vec{\rho}_{c}^{~\text{eq}}(\tau^{\prime})$ $=\int_{\tau_{0}}^{\tau}d\tau^{\prime}\partial_{\tau^{\prime}}\mathcal{W}_{\tau,\tau^{\prime}}\rho_{c}^{~\text{eq}}(\tau^{\prime})$
$=\mathcal{W}_{\tau,\tau^{\prime}}\rho_{c}^{~\text{eq}}(\tau^{\prime})\Bigg|_{\tau_{0}}^{\tau}-\int_{\tau_{0}}^{\tau}d\tau^{\prime}\mathcal{W}_{\tau,\tau^{\prime}}\partial_{\tau^{\prime}}\rho_{c}^{~\text{eq}}(\tau^{\prime})$
$=\mathcal{W}_{\tau,\tau^{\prime}}\rho_{c}^{~\text{eq}}(\tau^{\prime})\Bigg|_{\tau_{0}}^{\tau}-\int_{\tau_{0}}^{\tau}d\tau^{\prime}\mathcal{W}_{\tau,\tau^{\prime}}\hat{\boldsymbol{\nu}}_{c}(\tau^{\prime})\hat{\boldsymbol{\nu}}_{c}^{-1}(\tau^{\prime})\partial_{\tau^{\prime}}\rho_{c}^{~\text{eq}}(\tau^{\prime})$
$=\sum_{n=0}^{\infty}\mathcal{W}_{\tau,\tau^{\prime}}~\left[\hat{\boldsymbol{\nu}}_{c}^{-1}(\tau^{\prime})\partial_{\tau^{\prime}}\right]^{k}~\rho_{c}^{~\text{eq}}(\tau^{\prime})\Bigg|_{\tau_{0}}^{\tau}$
$=\hat{\mathcal{F}}^{-1}_{\tau,\tau_{0}}\sum_{n=0}^{\infty}\left[\hat{\boldsymbol{\nu}}^{-1}_{R}(\tau)\mathbf{\hat{D}}\right]^{k}~\rho^{~\text{eq}}(\tau)-\mathcal{W}_{\tau,\tau_{0}}~\sum_{n=0}^{\infty}\left[\hat{\boldsymbol{\nu}}^{-1}_{R}(\tau_{0})\mathbf{\hat{D}}\right]^{k}~\rho^{~\text{eq}}(\tau_{0})$ (233)
@@
$\vec{\rho}(\tau)=\hat{\mathcal{F}}_{\tau,\tau_{0}}\mathcal{W}_{\tau,\tau_{0}}\rho(\tau_{0})+\sum_{n=0}^{\infty}\left[\hat{\boldsymbol{\nu}}^{-1}_{R}(\tau)\mathbf{\hat{D}}\right]^{k}~\rho^{~\text{eq}}(\tau)-\hat{\mathcal{F}}_{\tau,\tau_{0}}\mathcal{W}_{\tau,\tau_{0}}~\sum_{n=0}^{\infty}\left[\hat{\boldsymbol{\nu}}^{-1}_{R}(\tau_{0})\mathbf{\hat{D}}\right]^{k}~\rho^{~\text{eq}}(\tau_{0})$ (234)
D.2 Free Streaming picture
@@
$\left[\partial_{\tau}+\hat{\mathbf{F}}_{S}\right]\vec{\rho}_{S}=\partial_{\tau}\vec{\rho}_{S}$ (240)
@@
$\hat{\mathcal{F}}_{\tau,\tau_{0}}=\mathcal{T}e^{-\int_{\tau_{0}}^{\tau}d\tau\hat{\textbf{F}}_{S}}$ (241)
@@
$\partial_{\tau}\hat{\mathcal{F}}_{\tau,\tau_{0}}=-\hat{\textbf{F}}_{S}~\hat{\mathcal{F}}_{\tau,\tau_{0}}$ (242)
@@
$\vec{\rho}^{\rm~G}(\tau)$ $=\int_{\tau_{0}}^{\tau}d\tau^{\prime}e^{-\xi^{\prime}}\hat{\mathcal{F}}_{\tau,\tau^{\prime}}\frac{\rho^{~\text{eq}}(\tau^{\prime})}{\tau_{R}}$ (246)
@@
$\vec{\rho}^{~\rm G}(\tau)$ $=\sum_{n=0}^{\infty}\frac{g_{n}(\tau,\tau_{0})}{n!}\sum_{k=0}^{n}h_{n,k}(\tau_{R},\mathbf{\hat{D}}\tau_{R})\left[-\tau_{R}\mathbf{\hat{D}}\right]^{k}\vec{\rho}^{~\text{eq}}$ (250)
$=\sum_{n=0}^{\infty}~\left(\sum_{k=n}^{\infty}\frac{g_{k}(\tau,\tau_{0})}{n!}h_{k,n}(\tau_{R},\mathbf{\hat{D}}\tau_{R})\right)~\left[-\tau_{R}\mathbf{\hat{D}}\right]^{n}\vec{\rho}^{~\text{eq}}$ (251)
D.3 Equilibrium gradients and closure
This section describe the standard formalism used to extract transport coefficients and higher order gradient expansion.
Consider the gradients of the equilibrium moments, $\left[-\hat{\boldsymbol{\nu}}^{-1}_{R}\mathbf{\hat{D}}\right]^{n}\vec{\rho}^{~\text{eq}}(\tau)$, where
$\mathbf{\hat{D}}\equiv\left[\partial_{\tau}+\hat{\mathbf{F}}_{\lambda}\nabla^{\lambda}+\hat{\mathbf{A}}\right$]\,. (252)
As $\vec{\rho}^{~\text{eq}}(\tau)$ is a function of $T,~\mu,~u^{\lambda}$, we can trade the space and time derivatives of $\vec{\rho}$ with space time derivatives of these hydrodynamic quantities.
$\partial_{\tau}\vec{\rho}^{~\text{eq}}$ $=\partial_{T}\vec{\rho}^{~\text{eq}}~\partial_{\tau}T+\partial_{\mu}\vec{\rho}^{~\text{eq}}~\partial_{\tau}\mu+\partial_{u^{\lambda}}\vec{\rho}^{~\text{eq}}~\partial_{\tau}u^{\lambda}$ (253)
$\nabla_{\sigma}\vec{\rho}^{~\text{eq}}$ $=\partial_{T}\vec{\rho}^{~\text{eq}}~\nabla_{\sigma}T+\partial_{\mu}\vec{\rho}^{~\text{eq}}~\nabla_{\sigma}\mu+\partial_{u^{\mu}}\vec{\rho}^{~\text{eq}}~\nabla_{\sigma}u^{\mu}$ (254)
We can now use the conservation equations,
$\partial_{\tau}n+n\,\partial_{\mu}u^{\mu}+\partial_{\mu}n^{\mu}$ $=0,$ (255)
$\partial_{\tau}\epsilon+(\epsilon+P+\Pi)\,\partial_{\mu}u^{\mu}-\pi^{\mu\nu}\,\nabla_{(\mu}u_{\nu)}$ $=0,$ (256)
$(\epsilon+P+\Pi)\,\partial_{\tau}u^{\alpha}-\nabla^{\alpha}(P+\Pi)+\Delta^{\alpha}{}_{\nu}\,\partial_{\mu}\pi^{\mu\nu}$ $=0.$ (257)
and the equation of state, $\epsilon=\epsilon(T,\mu),~n=n(T,\mu),~P=P(T,\mu)$ to eliminate the derivatives of $T$ and $\mu$ in favor of gradients of $u^{\lambda}$ and the anisotropic moments $\Pi$, $\pi^{\mu\nu}$. Now any gradient of $\Pi$ or $\pi^{\mu\nu}$ appear in the next order in expansion. These can be replaced with the expressions found in the lower order expansion. In the scheme presented here, there is a crucial difference that this includes the transient corrections.
This procedure reduces all order gradients in term of the gradients of $u^{\mu}$ and the other hydrodynamic quantities $\varepsilon,n,P,\Pi,\pi^{\mu\nu}$. In general we can write
$\sum_{k=0}^{n}\left[-\hat{\boldsymbol{\nu}}^{-1}_{R}\mathbf{\hat{D}}\right]^{k}\vec{\rho}^{~\text{eq}}(\tau)=\vec{\rm a}^{~(n)}~\epsilon+\vec{b}^{~(n)}~{\rm n}+\vec{\rm c}^{~(n)}~\Pi+\vec{d}_{\lambda\sigma}^{~(n)}~\pi^{\lambda\sigma}$ (258)
Where the coefficients are functions of $\varepsilon,n,P,\partial_{\lambda}u^{\sigma}$. Consider the projection of the above equation to the hydrodynamic subspace,
$\left(\vec{\rm a}^{~(n)}_{\mathcal{H}}~\varepsilon+\vec{b}^{~(n)}_{\mathcal{H}}~{\rm n}+\vec{\rm c}^{~(n)}_{\mathcal{H}}~\Pi+\vec{d}_{\mathcal{H}~\lambda\sigma}^{~(n)}~\pi^{\lambda\sigma}\right)_{i}$ $=(\vec{\rm a}^{~(n)}_{\mathcal{H}})_{i}\varepsilon+(\vec{b}^{~(n)}_{\mathcal{H}})_{i}~{\rm n}+(\vec{\rm c}^{~(n)}_{\mathcal{H}})_{i}~\Pi+(\vec{d}_{\mathcal{H}~\lambda\sigma}^{~(n)})_{i}~\pi^{\lambda\sigma}$
$=\sum_{h=0}^{N}(\mathbf{\hat{\Delta}}_{i,j})_{h}(\vec{\pi}_{\mathcal{H}})_{h}$ (259)
where $\mathbf{\hat{\Delta}}_{i,j}$ is an $N\times N$ matrix and we defined,
$\vec{\pi}_{\mathcal{H}}=\begin{bmatrix}\epsilon\\ \rm n\\ \Pi\\ \pi^{0,0}\\ \pi^{0,1}\\ \vdots\\ \pi^{4,4}\end{bmatrix}$ (260)
containing the fourteen independent hydrodynamic variables and
$\mathbf{\hat{\Delta}}=\begin{bmatrix}(\vec{a}){0}&(\vec{b}){0}&(\vec{c}){0}&(\vec{d}_{1,0})_{0}&\dots&(\vec{d}_{4,4})_{0}\\ (\vec{a}){1}&(\vec{b}){1}&(\vec{c}){1}&(\vec{d}_{1,0})_{1}&\dots&(\vec{d}_{4,4})_{1}\\ (\vec{a}){2}&(\vec{b}){2}&(\vec{c}){2}&(\vec{d}_{1,0})_{2}&\dots&(\vec{d}_{4,4})_{2}\\ (\vec{a}){3}&(\vec{b}){3}&(\vec{c}){3}&(\vec{d}_{1,0})_{3}&\dots&(\vec{d}_{4,4})_{3}\\ \vdots&\vdots&\vdots&\vdots&\dots&\vdots\\ (\vec{a}){14}&(\vec{b})_{14}&(\vec{c})_{14}&(\vec{d}_{1,0})_{14}&\dots&(\vec{d}_{4,4})_{14}\end{bmatrix}$ (261)
is a $14\times 14$ matrix corresponding to the the fourteen independent hydrodynamic variables.